How do you solve this question?:
The sum of two positive integers is 25. The differnece of their squares is 275. Find the intergers.
All i really need for this prolem is the equation using the quadraic equation.
Sort of set up like this
ex: x^2+2x=48
Something like that....

PLEASE HELP ME I AM DESPERATE!

Answer by mananth(16946)

The sum of two positive integers is 25. The differnece of their squares is 275. Find the intergers.
All i really need for this prolem is the equation using the quadraic equation.
Sort of set up like this
ex: x^2+2x=48
..
Let one number be x
the other number will be 25-x
..
The sum of their squares = 275
x^2 -( 25-x)^2= 275
x^2 -625+50x-x^2 =275
50x = 900
x=18
other number = 7
..
CHECK
18^2-7^2 =275